# Mixture of Scope Experts at Test: Generalizing Deeper Graph Neural Networks with Shallow Variants

### v3

@@ Appendix G Additional Experiments and Analysis
tuned-GAT [44] PathMLP [83] GNNDLD [6]
@@
NID [43] Dual-Net GNN [48] NHGCN [22]
@@
Polynormer [15] ACM-GCN++ [41] ACM-Snowball-3 [41]
@@ G.1 Leaderboard Comparison
Since Moscat is flexible for GNNs with various architectures, can GNN-Moscats achieve state-of-the-art performance? To answer this, Table 11 summarizes our comparison with the top 3 methods reported on amazon-ratings, Penn94, and PubMed leaderboards from Paper With Code. To the best of our knowledge, we find two methods, NID [43] and PathMLP [83], have reported top performance but have not been included in the leaderboard. We also include this method in our comparison. The result demonstrates the superior performance of GNN-Moscat. Note that we follow the standard setup for training the base GNN models and do not include any additional tricks.
@@ G.2 Runtime Analysis
Let $L^{\prime}$ denote the number of MLP layers and $F_{\text{hid}}$ denote the hidden dimension. The time complexity of $\mathcal{F}$ is bounded by $O\left(CL_{\text{max}}|{\mathcal{V}}|F_{\text{hid}}+|{\mathcal{V}}|L_{\text{max}}^{2}+L^{\prime}|{\mathcal{V}}|F_{\text{hid}}^{2}\right)$.
@@
    Moscat can be flexibly applied to shallow GNNs only while maintaining significant accuracy improvements, reducing the computational overhead of training deeper GNNs. As shown in Figure 6 and 7, the majority of accuracy gains from Moscat occur with GNNs of depth ${L\leq 4}$.
@@
G.3 Depth Analysis
We investigate how GNN-Moscat performance changes when varying the maximum depth $L_{\text{max}}$ of GNN experts. Figure 6 and Figure 7 compare the performance of GNN-Moscat and corresponding GNN on amazon-ratings dataset and arxiv-year dataset, respectively. As depth increases, GNN performance first increases and then decreases or remains unchanged, which is undesirable given deeper GNNs’ superior expressive power [13]. In contrast, GNN-Moscat leverages depth more effectively, showing a consistent upward trend in performance as expert depth increases. This indicates a promising characteristic of GNN-Moscat: it does not require careful tuning of $L_{\text{max}}$ tuning to achieve good performance. We found that $L_{\text{max}}=6$ is a good trade-off between accuracy and computational overhead for all datasets and all base GNNs. For cases with strict runtime requirements, setting $L_{\text{max}}=2$ for GNN-Moscat consistently achieves performance that is better than or comparable to the best GNN with depths between 1 and 10.
@@ G.3 Depth Analysis
Figure 8: Moscat performance upon classic GNNs (i.e., SGC and GCN) and GNNs with soft scoping methods (i.e., GAMLP and ACMGCN). GAMLP and ACMGCN adaptively learn the scope of SGC and GCN, respectively.
We further investigate how Moscat improves over two representative state-of-the-art deeper GNNs with soft scoping, ACMGCN [82] and GAMLP [41]. ACMGCN follows the message-passing architecture (e.g., GCN) while GAMLP follows the decoupled architecture (e.g., SGC). Both ACMGCN and GAMLP employ cross-layer gating to improve model performance when going deep. Figure 8 compares three model types at different depths: (1) classic GNNs (GCN and SGC), (2) deeper GNNs (ACMGCN and GAMLP), and (3) their corresponding Moscat variants. We observe that SGC and GCN performance often drops rapidly beyond 2 layers. However, with the gating mechanism, GAMLP and ACM-GCN show improved performance up to 4 layers and then exhibit flat curve behavior [21] (i.e., performance stabilizes with increasing depth). Moreover, GNNs with Moscat consistently perform best across all depths and show continuous accuracy improvements up to 6-10 layers, demonstrating the superiority of Moscat in leveraging deeper GNNs.
Figure 8 also reveals an interesting insight that Moscat particularly favors the message-passing architecture than the decoupled one for GNN experts. As depth increases from 1 to 10 layers, message-passing GNNs with Moscat achieve 4$\sim$10accuracy improvements (Figure 8 top), while decoupled GNNs with Moscat show only 1$\sim$2improvements (Figure 8 bottom). This difference likely stems from message-passing GNNs’ superior expressivity, enabling GNNs of different depths to generate more diverse predictions.
@@
G.4 Empirical Evidence of Deeper Soft Scoping Methods Suffer from Overfitting
We plot the training and test accuracy of a representative soft scoping method ACMGCN [41] (see Figure 9). The figure demonstrates that the 6-layer ACMGCN achieves higher training accuracy than the 2-layer version. However, in homophily regions, the 6-layer model exhibits lower test accuracy. This suggests that while deeper soft scoping methods offer greater expressive power, they are also prone to overfitting. Considering the architecture of soft scoping methods, we speculate this issue arises because shallow layers/experts in soft scoping models may overfit to noise from higher-hop neighbors.
G.5 Supplementary Details for Heterophily-Biased Sample Filtering
@@
G.6 How Expert Training Affects Moscat Performance
@@
Table 15: Test accuracy comparisons across different hidden dimensions for GNN experts. We set $L_{\text{max}}=6$ and fix other GNNs and Moscat hyperparameters. For GCN and GAT, we report their best accuracy across configurations with 1-6 layers. ($0$$\pm$0.00) denotes the corresponding training accuracy.
@@
Table 16: Test accuracy comparisons across different training epochs for GNN experts. We select checkpoints at 300, 1000, and 2000 training epochs for each expert, representing underfitting, well-fitting, and overfitting states, respectively. We set $L_{\text{max}}=6$, expert droput to 0, and fix other GNNs and Moscat hyperparameters. For GCN and GAT, we report their best accuracy across configurations with 1-6 layers. ($0$$\pm$0.00) denotes the corresponding training accuracy.
@@ G.7 Interpretation and Visualization
To analyze the behavior of AS, we use t-SNE [60] to visualize the hidden embeddings from three scope experts, which are GAT models with layers ranging from 0 to 2, on the Penn94 dataset (Figure 12 Left). Specifically, we add an output linear layer to each GAT model and visualize the hidden embeddings just before this layer. Since AS predicts the optimal scope expert (i.e., Scope-0, Scope-1, or Scope-2) for each node, we highlight nodes according to the predicted scope expert in the embedding visualization (Figure 12 Right).
@@
G.8 More Empirical Findings of Performance Disparity across Scope Experts
@@ G.8.1 Overlapping Ratio
    Scope-0 and Scope-1 experts diverge most: Across architectures and datasets, experts with scopes of 0 and 1 exhibit the lowest overlap with other experts, suggesting they capture unique information relative to larger-scope variants.
@@
    Dataset homophily drives overlap: Heterophilous datasets (e.g., Chameleon, Squirrel, and snap-patents with average node homophily less than 0.2) show consistently low overlaps (often $<$ 0.7), whereas homophilous datasets (e.g., arxiv-year, PubMed) yield high overlaps (often $>$ 0.7).
@@
    Architecture-dependent diversity: Attention-based models (GAT, ACMGCN) produce lower overlaps and more varied patterns, particularly GAT on Penn94 and ACMGCN on Pubmed, indicating greater specialization among experts. In contrast, the decoupled SGC architecture exhibits uniformly higher overlaps, reflecting more redundant predictions among its experts.
@@ G.8.2 Performance Disparity
G.8.3 Ensembling Upper-bound
@@
G.9 Analysis of the Gating Weights
Figure 18 and Figure 19 illustrate the distributions of gating weights for various experts in Moscat. The results indicate that Moscat effectively learns gating weights, thereby preventing the collapse issue (i.e., some experts receive extremely small weights across nodes) as noted in prior studies [56, 59]. Moreover, the gating model tends to assign larger weights with greater variance to shallow experts (e.g., Scope-0 and Scope-1) compared to deeper ones. This behavior suggests that Moscat learns more personalized weights for shallow experts, while it adopts a more uniform weighting scheme for deeper experts. One plausible explanation for this trend is that increasing the scope size leads to a higher likelihood of overlapping neighbors among different nodes. Therefore, the gating model tends to learn the same weight on deeper experts for these nodes.

### v4

@@ Appendix G Additional Experiments and Analysis
tuned-GAT [tuned] PathMLP [pathmlp] GNNDLD [gndld]
@@
NID [nid] Dual-Net GNN [dualnet] NHGCN [nhgcn]
@@
Polynormer [polynormer] ACM-GCN++ [acmgcn] ACM-Snowball-3 [acmgcn]
@@ G.1 Leaderboard Comparison
Since Moscat is flexible for GNNs with various architectures, can GNN-Moscats achieve state-of-the-art performance? To answer this, Table 11 summarizes our comparison with the top 3 methods reported on amazon-ratings, Penn94, and PubMed leaderboards from Paper With Code. To the best of our knowledge, we find two methods, NID [nid] and PathMLP [pathmlp], have reported top performance but have not been included in the leaderboard. We also include this method in our comparison. The result demonstrates the superior performance of GNN-Moscat. Note that we follow the standard setup for training the base GNN models and do not include any additional tricks.
@@ G.2 Runtime Analysis
Let $L^{\prime}$ denote the number of MLP layers and $F_{\text{hid}}$ denote the hidden dimension. The time complexity of the gating model $\mathcal{F}$ is bounded by $O\left(CL_{\text{max}}|{\mathcal{V}}|F_{\text{hid}}+|{\mathcal{V}}|L_{\text{max}}^{2}+L^{\prime}|{\mathcal{V}}|F_{\text{hid}}^{2}\right)$.
@@
    Moscat can be flexibly applied to shallow GNNs only while maintaining significant accuracy improvements, reducing the computational overhead of training deeper GNNs. As shown in Figure 7 and 8, the majority of accuracy gains from Moscat occur with GNNs of depth $L\leq 4$}.
@@
G.3 Space Analysis
Moscat allows each expert and the gating model to be trained separately. Practitioners can flexibly adjust the parallelism of expert training to balance runtime and memory consumption. Let $L_{\max}$ denote the maximum depth of GNN experts, $F_{\text{hid}}$ denote the hidden dimension, and $\mathcal{V}_{\exp}$ and $\mathcal{V}_{\text{gate}}$ denote the training sets for experts and the gating model, respectively. Taking GCN as an example, we have two boundary cases:
  • If we prioritize minimal memory overhead, we can train each expert sequentially. The space complexity is therefore bounded by training the deepest GNN expert: $O\!\left(L_{\max}|\mathcal{V}_{\exp}|F_{\text{hid}}+L_{\max}F_{\text{hid}}^{2}\right).$
  • If we prioritize minimal runtime, we can train all experts in parallel. The space complexity then becomes proportional to the total number of layers $(1+2+\cdots+L_{\max})$ across all experts: $O\!\left(L_{\max}^{2}|\mathcal{V}_{\exp}|F_{\text{hid}}+L_{\max}^{2}F_{\text{hid}}^{2}\right).$
The memory consumption of the gating model training is relatively small and won’t become a bottleneck: $O\!\left(L_{\max}|\mathcal{V}_{\text{gate}}|F_{\text{hid}}+L_{\max}F_{\text{hid}}^{2}\right)$, since $F_{\text{hid}}\ll|\mathcal{V}_{\text{gate}}|\ll|\mathcal{V}_{\exp}|$. In conclusion, with comparable runtime, Moscat requires approximately $L_{\max}$ times more memory than a single $L_{\max}$-layer GNN during training.
G.4 Depth Analysis
We investigate how GNN-Moscat performance changes when varying the maximum depth $L_{\text{max}}$ of GNN experts. Figure 7 and Figure 8 compare the performance of GNN-Moscat and corresponding GNN on amazon-ratings dataset and arxiv-year dataset, respectively. As depth increases, GNN performance first increases and then decreases or remains unchanged, which is undesirable given deeper GNNs’ superior expressive power [provable]. In contrast, GNN-Moscat leverages depth more effectively, showing a consistent upward trend in performance as expert depth increases. This indicates a promising characteristic of GNN-Moscat: it does not require careful tuning of $L_{\text{max}}$ tuning to achieve good performance. We found that $L_{\text{max}}=6$ is a good trade-off between accuracy and computational overhead for all datasets and all base GNNs. For cases with strict runtime requirements, setting $L_{\text{max}}=2$ for GNN-Moscat consistently achieves performance that is better than or comparable to the best GNN with depths between 1 and 10.
@@ G.4 Depth Analysis
G.5 Empirical Evidence of Deeper Soft Scoping Methods Suffer from Overfitting
We plot the training and test accuracy of a representative soft scoping method ACMGCN [acmgcn] (see Figure 9). The figure demonstrates that the 6-layer ACMGCN achieves higher training accuracy than the 2-layer version. However, in homophily regions, the 6-layer model exhibits lower test accuracy. This suggests that while deeper soft scoping methods offer greater expressive power, they are also prone to overfitting. Considering the architecture of soft scoping methods, we speculate this issue arises because shallow layers/experts in soft scoping models may overfit to noise from higher-hop neighbors.
G.6 Supplementary Details for Heterophily-Biased Sample Filtering
@@
G.7 How Expert Training Affects Moscat Performance
@@
Table 15: Test accuracy comparisons across different hidden dimensions for GNN experts. We set $L_{\text{max}}=6$ and fix other GNNs and Moscat hyperparameters. For GCN and GAT, we report their best accuracy across configurations with 1-6 layers. (00.00 $\pm$0.00) denotes the corresponding training accuracy.
@@
Table 16: Test accuracy comparisons across different training epochs for GNN experts. We select checkpoints at 300, 1000, and 2000 training epochs for each expert, representing underfitting, well-fitting, and overfitting states, respectively. We set $L_{\text{max}}=6$, expert droput to 0, and fix other GNNs and Moscat hyperparameters. For GCN and GAT, we report their best accuracy across configurations with 1-6 layers. (00.00 $\pm$0.00) denotes the corresponding training accuracy.
@@ G.8 Interpretation and Visualization
To analyze the behavior of AS, we use t-SNE [tsne] to visualize the hidden embeddings from three scope experts, which are GAT models with layers ranging from 0 to 2, on the Penn94 dataset (Figure 12 Left). Specifically, we add an output linear layer to each GAT model and visualize the hidden embeddings just before this layer. Since AS predicts the optimal scope expert (i.e., Scope-0, Scope-1, or Scope-2) for each node, we highlight nodes according to the predicted scope expert in the embedding visualization (Figure 12 Right).
@@
G.9 More Empirical Findings of Performance Disparity across Scope Experts
@@ G.9.1 Overlapping Ratio
    Scope‑0 and Scope‑1 experts diverge most: Across architectures and datasets, experts with scopes of 0 and 1 exhibit the lowest overlap with other experts, suggesting they capture unique information relative to larger‑scope variants.
@@
    Dataset homophily drives overlap: Heterophilous datasets (e.g., Chameleon, Squirrel, and snap-patents with average node homophily less than 0.2) show consistently low overlaps (often << 0.7), whereas homophilous datasets (e.g., arxiv-year, PubMed) yield high overlaps (often >> 0.7).
@@
    Architecture‑dependent diversity: Attention‑based models (GAT, ACMGCN) produce lower overlaps and more varied patterns, particularly GAT on Penn94 and ACMGCN on Pubmed, indicating greater specialization among experts. In contrast, the decoupled SGC architecture exhibits uniformly higher overlaps, reflecting more redundant predictions among its experts.
@@ G.9.2 Performance Disparity
G.9.3 Ensembling Upper-bound
@@
G.10 Analysis of the Gating Weights
Figure 18 and Figure 19 illustrate the distributions of gating weights for various experts in Moscat. The results indicate that Moscat effectively learns gating weights, thereby preventing the collapse issue (i.e., some experts receive extremely small weights across nodes) as noted in prior studies [graphgps, shazeer2017outrageously]. Moreover, the gating model tends to assign larger weights with greater variance to shallow experts (e.g., Scope-0 and Scope-1) compared to deeper ones. This behavior suggests that Moscat learns more personalized weights for shallow experts, while it adopts a more uniform weighting scheme for deeper experts. One plausible explanation for this trend is that increasing the scope size leads to a higher likelihood of overlapping neighbors among different nodes. Therefore, the gating model tends to learn the same weight on deeper experts for these nodes.